OpenStudy (anonymous):

find the standard deviation for the following set of info; 1 1 1 1 1 1 1 1

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

wouldn't it be zero since the set is all the same numbers

OpenStudy (anonymous):

i guess but how do you find standard deviation ex 3 5 5 7 7 9

OpenStudy (anonymous):

it's a little confusing you take the square of all the numbers in the set and find the mean of that and you subtract it by the normal mean of the set squared and you take the square root of all of that eg. [(9+25+25+49+49+81)/6] - (3+5+5+7+7+9/6)^2 and whatever you get you take the square root

OpenStudy (amistre64):

thats not how i learnt it ....

OpenStudy (anonymous):

it's been awhile i just googled it if you have a better way to explain it then go ahead lol

OpenStudy (anonymous):

i'm basically taking the square root of var(x)

OpenStudy (amistre64):

find the mean of the set find the difference between the mean and each datapoint square those differences and add em up \[\sum (x-\mu)^2\] divide by n (for population) ; or n-1 (for samples) \[\frac{\sum (x-\mu)^2}{N,(n-1)}\] this is variance; standard deviation is just the sqrt of variance

OpenStudy (anonymous):

to find the mean do i square the numbers

OpenStudy (amistre64):

no, to find the mean you add the numbers and divide by how may there are

OpenStudy (amistre64):

\[mean(a_1,a_2,a_3,...,a_n)=\frac{a_1+a_2+a_3+...+a_n}{n}\]

OpenStudy (anonymous):

ok then what

OpenStudy (amistre64):

go to the next step ...

OpenStudy (amistre64):

find the difference between the mean and the data points

OpenStudy (anonymous):

i dont understand data point

OpenStudy (amistre64):

the numbers you are given are the data results of the tests/experiments/goings-on

OpenStudy (amistre64):

you are given a set of numbers, of data points, to manuver with

OpenStudy (anonymous):

can you show me

OpenStudy (amistre64):

can i show you what your numbers are?

OpenStudy (anonymous):

3 5 5 5 7 7 8

OpenStudy (amistre64):

ok, and the mean of those is?

OpenStudy (anonymous):

opps take off one 5 and replace 8 wiith 9

OpenStudy (anonymous):

ok 6

OpenStudy (amistre64):

subtract the mean, you say its 6, from each number and rewrite the set

OpenStudy (amistre64):

it doesnt matter if you subtract the mean from the number or the number from the mean; its the difference we want; sign doesnt matter

OpenStudy (anonymous):

-3 -1 -1 1 1 3 ok

OpenStudy (amistre64):

those look fine now square them

OpenStudy (anonymous):

9 1 1 1 1 9

OpenStudy (amistre64):

good, now add those all together

OpenStudy (anonymous):

22

OpenStudy (amistre64):

almost there; do you know if this data represents a sample of a population?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

well, when you know if its a pop. (divide by the number of data points) or a samp (number of data points minus 1) you have 6 numbers; so lets assume its a population just for closure :) 22/6 = variance sqrt(22/6) = standard deviation

OpenStudy (amistre64):

if it was a sample; wed of divided by 5 in this case

OpenStudy (anonymous):

ok thnks so much!!!!!!!:):) u rock

OpenStudy (amistre64):

oure welcome, hope it helped :)